Calculate the molecular masses of the following molecules.

MM (g/mol) of PBr3?
MM (g/mol) of WO3?
MM (g/mol) of C7H16?
MM (g/mol) of Ca(H2PO4)2
1
Expert's answer
2020-11-14T13:55:14-0500

MM (g/mol) of PBr3 "= 30.97 +(3\times 79.9) = 270.67"

MM (g/mol) of WO3 "= 183.85 + (3 \times 15.99) = 231.82"

MM (g/mol) of C7H16 "= (7 \times 12.01) + (16 \times 1) = 100.07"

MM (g/mol) of Ca(H2PO4)2 "= 40.08 + (2 \times 2 \times 1) + (2 \times 30.97) + (2 \times 4 \times 15.99) = 40.08 + 4 + 61.94 + 127.92 = 233.94"